# Requirements Coverage-Guided Minimization for Natural Language Test Cases

### v2

@@ 4.2. Dataset
For example, for a test suite $\mathcal{T}$ with three test cases that detect $3$, 22, and $3$ faults, respectively. The total number of unique faults detected by $\mathcal{T}$ is $4$, the redundancy level of $\mathcal{T}$ is therefore $(3+2+3)/4=2$. Similarly, the redundancy level of our dataset is $2610/220=11.86$.
@@ 4.3. Text Representations
Inspired by prior work on vector representations of requirements (zhao2025machine; abbas2025requirements) and test cases (viggiato2022identifying), we employ a variety of text embedding techniques to convert test cases into vector representations. Specifically, we use TF-IDF, USE, Amazon Titan Text Embedding V2, and LongT5 to generate sentence-level embeddings, while Word2Vec, GloVe, and FastText are used to obtain word-level embeddings. Due to limitations in computational resources, data privacy constraints, and restrictions on input token length, we were unable to experiment with models such as BERT (devlin2019bert) and Sentence-BERT (DBLP:conf/emnlp/ReimersG19).
@@
TF-IDF. TF-IDF has demonstrated promising performance in identifying similar test cases expressed in natural language (viggiato2022identifying). We also employed TF-IDF to extract numerical vector representations of the test cases. For each word, we calculated its importance within a single test case relative to its occurrence across all other test cases.
@@
USE (Universal Sentence Encoder) (DBLP:journals/corr/abs-1803-11175). USE encodes text into high-dimensional vectors, capturing semantic information at the sentence level, making it particularly effective for tasks such as semantic similarity, text classification, clustering, and information retrieval (DBLP:journals/corr/abs-1803-11175).
@@
LongT5 (DBLP:conf/naacl/GuoAUONSY22). LongT5 is an extension of the T5 architecture designed to efficiently handle long sequences (DBLP:conf/naacl/GuoAUONSY22). It incorporates sparse attention and reversible layers to reduce memory usage, making it well-suited for tasks like long-document summarization and retrieval-augmented generation. We selected the local attention mechanism for LongT5 because it yielded a higher FDR for TSM than the global attention mechanism.
@@
Amazon Titan Text Embedding V2 (aws2024titantv2). For LLM-based embeddings, we use Amazon Titan Text Embedding V2, provided by our industrial partners. The model takes test case steps as input and produces a $1,024$-dimensional vector representation for test cases.
@@
Word2Vec (DBLP:journals/corr/abs-1301-3781). Word2Vec has been shown to be effective for identifying similar test cases (viggiato2022identifying). In this study, we used the Continuous Bag-of-Words (CBOW) architecture with a window size of $10$ for Word2Vec, as this configuration yielded the highest FDR for TSM. This was determined through experiments evaluating both the CBOW and Skip-Gram architectures with window sizes ranging from $2$ to $10$, in increments of $2$. We trained Word2Vec models on the entire test case corpus and transformed each word into $300$-dimensional numerical vectors.
@@
GloVe (DBLP:conf/emnlp/PenningtonSM14). GloVe is an unsupervised learning algorithm that leverages global word-word co-occurrence statistics from a corpus to learn word embeddings (DBLP:conf/emnlp/PenningtonSM14). Its core idea is that the ratio of co-occurrence probabilities can capture meaningful linear substructures within the word vector space. We selected a window size of $2$ for GloVe based on its performance in hyperparameter tuning experiments with window sizes ranging from $2$ to $10$, in increments of $2$. As in Word2Vec, we used GloVe to generate a $300$-dimensional numerical vector representation for each word in the test cases.
@@
FastText (DBLP:journals/tacl/BojanowskiGJM17). FastText is a word embedding model proposed by Bojanowski et al. (DBLP:journals/tacl/BojanowskiGJM17), which extends Word2Vec by incorporating subword information. Each word is represented as a combination of character n-grams, enabling the model to better handle rare and out-of-vocabulary words, and to generalize more effectively—particularly in morphologically rich or noisy text. We used the CBOW architecture with a window size of $8$ as the FastText configuration, based on experiments evaluating both CBOW and Skip-Gram architectures with window sizes ranging from $2$ to $10$ in increments of $2$. We used FastText to convert each word in the test cases into a $300$-dimensional numerical vector.
4.4. Baselines
@@
We compare RTM against (1) Random Minimization, a standard TSM baseline (pan2023atm), (2) FAST-R approaches (cruciani2019scalable), a set of novel TSM approaches with two versions: fixed-size budget version and adequate version. The former satisfies the pre-defined minimization budget constraint, whereas the latter satisfies the full requirement coverage constraint. While being a widely used approach for TSM (khan2018systematic), Greedy-based TSM approaches iteratively select the test cases that maximize the coverage until the full coverage constraint is satisfied (miranda2017scope; noemmer2019evaluation). However, in our dataset, each test case covers only one requirement, rendering this option equivalent to Random Minimization with a requirement coverage constraint, as described below.
@@ 4.4.1. Random Minimization
Since the goal of TSM is to reduce the number of test cases and, consequently, lower testing costs, the most straightforward approach is to randomly remove test cases to meet the minimization budget (DBLP:journals/tse/AliBHP10). We evaluate the performance of RTM against Random Minimization with and without the requirement coverage constraint, denoted as RM-Req and RM-NoReq, respectively. For the version with the constraint, we applied the same selection strategy as GA initialization described in Section 3.4.1. For the version without the constraint, we randomly select $k$ test cases from the test suite, ensuring that $k$ adheres to the minimization budget.
4.4.2. FAST-R (cruciani2019scalable)
FAST-R is a family of similarity-based TSM approaches, including FAST++, FAST-CS, FAST-pw, and FAST-all. FAST++ and FAST-CS utilize k-means++ clustering (DBLP:conf/soda/ArthurV07) and constructed coresets (DBLP:conf/kdd/BachemL018). In contrast, FAST-pw and FAST-all leverage minhashing and locality-sensitive hashing (DBLP:books/cu/LeskovecRU14). FAST-R can be applied in two scenarios: the fixed-budget and adequate scenarios. The fixed-budget version of FAST-R minimizes the test suite within a predefined budget, while the adequate version prioritizes preserving the requirement coverage of the minimized test suite over adhering to the budget. One limitation of FAST-R is that it cannot simultaneously guarantee both a minimization budget and full requirement coverage. We compare RTM against FAST-R under both scenarios where FAST-R is applicable (i.e., fixed-budget and adequate test suite minimization). For FAST-R, we relied on the publicly available replication package [^3] https://github.com/ICSE19-FAST-R/FAST-R provided by its authors to ensure a fair and consistent comparison.
@@ 5.1. Performance of RTM under different configurations (RQ1)
Interestingly, despite being a simple, sparse representation, TF-IDF outperformed all other sentence-level embeddings. The characteristics of our industrial test cases can explain this. As shown in Figure 1, the test cases in our dataset follow structured and repetitive templates with little linguistic variation. The steps in the test cases differ only in variable names or parameter values, and TF-IDF precisely captures these differences through its word-frequency mechanism, down-weighting common words that appear in every test case and up-weighting rare words, including distinct variable names and parameter values. In contrast, other sentence-level embeddings, while capable of capturing deep semantic relationships, tend to smooth out subtle lexical differences that are crucial for distinguishing test cases in our dataset. This suggests that in template-based, domain-specific test descriptions, TF–IDF, which suppresses common terms and emphasizes rare tokens, often outperforms semantically rich embeddings that smooth over token-level distinctions.
@@
Answering RQ1: TF-IDF with cosine similarity yields the highest overall FDR ($86.09\%$) under Preprocessing Method $PM3$ and Initialization Strategy 2 at the $50\%$ minimization budget.
Table 1. FDR (%) across different configurations (50% minimization budget)
@@ 5.2. Performance of RTM compared with baselines (RQ2)
We evaluated RTM against Random Minimization, which also employs Initialization Strategy 2, two versions of FAST-R (cruciani2019scalable)—the fixed-budget and adequate versions. For the fixed-budget version of FAST-R, we varied the minimization budget from $10\%$ to $90\%$, in increments of $10\%$. The fixed-budget version of FAST-R and RM-NoReq do not guarantee the 100% requirement coverage and are included in this RQ solely for FDR comparison. In the adequate scenario, we first ran FAST-R to determine the resulting minimization budget. This budget was then applied consistently to RTM and other baselines for experimental comparison. Notably, due to redundancy in our dataset, as few as $7\%$ (i.e., $54$ test cases, one test case per requirement) are sufficient to achieve $100\%$ requirement coverage.
@@
We observe that RTM consistently outperforms Random Minimization and FAST-R approaches in terms of FDR across various minimization budgets. The performance gap is particularly notable at mid-range budgets ($30\%$ to $60\%$). For instance, at a $40\%$ minimization budget, RTM achieves an FDR of $78.7\%$, whereas RM-Req and FAST++ yield approximately $65.5\%$ and $56.5\%$, respectively. This demonstrates RTM’s superior ability to retain fault-revealing test cases, especially under mid-range budget constraints. At lower and higher minimization budgets, the FDR of all techniques tends to converge toward either 0 or $1$, thus making the difference in FDR between RTM and other techniques smaller. However, mid-range minimization budgets are more important in practice, as they offer more useful trade-offs between cost savings and fault detection.
@@
Table 2. FDR (%) and Requirement Coverage (%) Comparison at Adequate Scenario for FAST-R
@@
Figure 3. Comparison of FDR across minimization budgets for RTM, FAST-R, and Random Minimization, alongside the theoretical upper bound.:
@@ 5.3. Impact of test suites redundancy (RQ3)
As expected, the FDR of all techniques improves as the test suite redundancy level increases, indicating that the TSM techniques tend to perform better–achieving higher FDR–when applied to more redundant test suites. This trend, however, becomes less evident at higher budgets ($80\%$ and $90\%$), where the FDR of all techniques converge toward $1$. Random Minimization (RM-Req and RM-NoReq) slightly outperforms other FAST-R variants but consistently falls short of RTM. We also observed that across all minimization budgets and redundancy levels, RM-Req consistently outperformed RM-NoReq in FDR. This demonstrates that, for our dataset, explicitly ensuring that every requirement is covered not only preserves full requirement coverage but also significantly boosts the test suite’s ability to detect faults. In other words, prioritizing requirement coverage is an effective heuristic for improving fault detection under budgetary constraints.
@@
Figure 4. Comparison of FDR Across Varying Redundancy Levels for RTM and Baseline Approaches Under Seven Minimization Budgets:
@@ 5.4. Scalability of RTM compared with baselines (RQ4)
We observe that, although being slower than all other baselines, RTM minimizes test cases in seconds–$1.32$s to $4.44$s in terms of total minimization time–for test suite size from $187$ to $616$, which is computationally efficient. Moreover, the fact that the total minimization time of RTM follows a linear relationship with the test suite size indicates that RTM is scalable for larger test suites. According to the fitted linear regression model, when the test suite size $n$ is 200,000200{,}000, the predicted total minimization time of RTM is $6.217686\times 10^{-3}\,n+2.523669\approx 1246.06$ seconds, whereas FAST-CS (the fastest FAST-R approach) is $1.195291\times 10^{-4}\,n+0.051979\approx 23.96$ seconds and RM-Req is $9.062634\times 10^{-7}*n+0.000472\approx 0.18$ seconds. Thus, the time differences are $20.37$ mins and $20.76$ mins when compared to FAST-CS and RM-Req, respectively. Given the fact that, unlike test case selection and prioritization, test suite minimization only performs occasionally, such as major releases with many new added test cases and redundant old test cases (DBLP:journals/tse/PanGB24; pan2023atm), such differences are practically acceptable.
@@
In summary, RTM runs within $5$ seconds across all evaluated test suites. Although being slower than the baselines, RTM scales approximately linearly with test suite sizes, and even for substantially larger suites, the absolute runtime gaps remain acceptable in practice—they are not expected to have practical consequences for TSM. Given that RTM consistently achieves much higher FDR across various datasets and minimization budgets than the baselines while preserving full requirement coverage, it is therefore a better option in many practical contexts.
@@
Figure 5. Scatter plots of the number of test cases and preparation time, search time and total minimization time (in sec) for RTM and baselines under 50% minimization budget:
@@ 7. Related Work
Several approaches have been proposed to support TSM, which can be classified in to three categories (khan2018systematic; cruciani2019scalable; pan2023atm): greedy (noemmer2019evaluation; miranda2017scope), clustering (coviello2018clustering; liu2011user; viggiato2022identifying; cruciani2019scalable), and search-based (zhang2019uncertainty; hemmati2013achieving; pan2023atm; DBLP:journals/tse/PanGB24; marchetto2017combining), plus hybrid combinations (xia2021test; anwar2019hybrid; yoo2010using) thereof.
@@
Greedy-based approaches follow a heuristic, step-by-step selection strategy, where the test case that provides the most benefit toward the objective is chosen at each iteration, until the given constraints (e.g., full requirement coverage or code path coverage) are satisfied. Miranda et al. (miranda2017scope) and Noemmer et al. (noemmer2019evaluation) both employed greedy heuristics for TSM based on statement coverage. Their approaches iteratively selected test cases to maximize code coverage, achieving reasonable FDRs while significantly reducing the number of executed test cases. Greedy-based approaches are computationally efficient and easy to implement, but may suffer from suboptimal global performance due to their locally optimal decisions. Moreover, given that, in our dataset, each test case covers only one requirement, the coverage-based greedy strategy–iteratively selecting the test cases that maximize the coverage until the full coverage constraint is satisfied–is equivalent to Random Minimization with the requirement constraint (RM-Req), as described in Section 4.4
@@
Clustering-based approaches group similar test cases based on syntactic or semantic similarities, such as those derived from natural language descriptions or execution behavior. The goal is to reduce redundancy by selecting representative test cases from each cluster. These approaches are particularly useful in settings where test cases are written in natural language and code coverage information is unavailable. Viggiato et al. (viggiato2022identifying) proposed a comprehensive approach that combines text embedding, similarity measurement, and clustering to identify and remove similar test cases written in natural language. They split the requirement test cases into test steps and then calculate the similarity between test steps. Cruciani et al. (cruciani2019scalable) proposed FAST-R, a black-box TSM approach that leverages the source code of test cases without requiring execution or coverage information. FAST-R transforms test code into vector representations using a term frequency model, followed by dimensionality reduction via random projection. Similarly, Coviello et al. (coviello2018clustering) proposed a clustering-based approach using Hierarchical Agglomerative Clustering to perform TSM.
@@
Search-based approaches formulate the minimization problem as an optimization task and utilize metaheuristic algorithms (e.g., GA, simulated annealing) to explore the space of possible test subsets. These methods can balance multiple conflicting objectives, such as minimizing the number of test cases while maximizing fault detection capability and maintaining requirement coverage. Hemmati et al. (hemmati2013achieving) applied a search algorithm that minimizes test suites based on model-derived test case similarity. Similarly, Zhang et al.(zhang2019uncertainty) proposed UncerTest, a model-based test case minimization framework that leverages multi-objective search algorithms, including NSGA-II, MOCell(nebro2009mocell), and SPEA2 (zitzler2001spea2). ATM (pan2023atm) is a black-box TSM approach that combines test code similarity analysis with evolutionary search. It represents test code using Abstract Syntax Trees and employs GA and Non-Dominated Sorting Genetic Algorithm II (meyarivan2002fast) as the underlying evolutionary algorithms. Furthermore, LTM (DBLP:journals/tse/PanGB24) leverages LLMs (CodeBERT, UniXcoder, and CodeLlama) to compute similarity between test cases and employs a GA to search for an optimal subset of test cases under a fixed budget. MORE+ (marchetto2017combining) is a multi-objective TSM approach that uses NSGA-II to optimize structural, functional, and cost-related objectives, aiming to enhance fault detection while reducing redundancy and execution time. Although typically more computationally intensive than greedy and clustering approaches, search-based techniques support multiple minimization objectives and often achieve better performance for TSM.
@@
Hybrid approaches integrate complementary methods into a unified framework to balance multiple objectives, including reducing test suite size, preserving fault detection capability, and satisfying requirement or coverage constraints. Xia et al. (xia2021test) apply clustering-based test suite reduction combined with evolutionary multi-objective optimization. Anwar et al. (anwar2019hybrid) combined a GA and particle swarm optimization to optimize regression test suites. Yoo et al. (yoo2010using) employ a hybrid multi-objective GA that integrates the efficient approximation of the greedy approach with the population-based GA’s ability to generate higher-quality Pareto fronts.
@@
While promising, existing techniques primarily focus on code-based solutions for minimization. In contrast, our work addresses minimization in the context of requirement-driven testing, where test cases are specified in natural language, which introduces distinct challenges. Moreover, current approaches do not explicitly support TSM under a fixed budget while simultaneously enforcing full requirement coverage as a hard constraint. To this end, we adopt a GA, as the SOTA TSM work (i.e., ATM and LTM), and augment it with an explicit requirement coverage constraint that guarantees $100\%$ coverage under a fixed minimization budget. Note that we did not use NSGA-II, as reported in the ATM work, since it achieves comparable performance while being substantially more time-consuming than the GA. To the best of our knowledge, this is also the first work to investigate the impact of redundancy levels on fault detection capability under such constraints, providing valuable insights for future research and practical applications. Note that we retained FAST-R as one of the baselines, though it cannot simultaneously satisfy both the fixed-size minimization budget and the full coverage adequacy constraint. This is because FAST-R can be applied in two scenarios (i.e., fixed-size budget and adequate), which enables experimental comparison with our approach.

### v3

@@ 4.2. Dataset
For example, for a test suite $\mathcal{T}$ with three test cases that detect 33, 22, and 33 faults, respectively. The total number of unique faults detected by $\mathcal{T}$ is 44, the redundancy level of $\mathcal{T}$ is therefore $(3+2+3)/4=2$. Similarly, the redundancy level of our dataset is $2610/220=11.86$.
@@ 4.3. Text Representations
Inspired by prior work on vector representations of requirements (Zhao and Alhoshan, 2025; Abbas et al., 2025) and test cases (Viggiato et al., 2022), we employ a variety of text embedding techniques to convert test cases into vector representations. Specifically, we use TF-IDF, USE, Amazon Titan Text Embedding V2, and LongT5 to generate sentence-level embeddings, while Word2Vec, GloVe, and FastText are used to obtain word-level embeddings. Due to limitations in computational resources, data privacy constraints, and restrictions on input token length, we were unable to experiment with models such as BERT (Devlin et al., 2019) and Sentence-BERT (Reimers and Gurevych, 2019).
@@
TF-IDF (Salton and Buckley, 1988). TF-IDF has demonstrated promising performance in identifying similar test cases expressed in natural language (Viggiato et al., 2022). We also employed TF-IDF to extract numerical vector representations of the test cases. For each word, we calculated its importance within a single test case relative to its occurrence across all other test cases.
@@
USE (Universal Sentence Encoder) (Cer et al., 2018). USE encodes text into high-dimensional vectors, capturing semantic information at the sentence level, making it particularly effective for tasks such as semantic similarity, text classification, clustering, and information retrieval (Cer et al., 2018).
@@
LongT5 (Guo et al., 2022). LongT5 is an extension of the T5 architecture designed to efficiently handle long sequences (Guo et al., 2022). It incorporates sparse attention and reversible layers to reduce memory usage, making it well-suited for tasks like long-document summarization and retrieval-augmented generation. We selected the local attention mechanism for LongT5 because it yielded a higher FDR for TSM than the global attention mechanism.
@@
Amazon Titan Text Embedding V2 (AWS, 2024). For LLM-based embeddings, we use Amazon Titan Text Embedding V2, provided by our industrial partners. The model takes test case steps as input and produces a $1,024$-dimensional vector representation for test cases.
@@
Word2Vec (Mikolov et al., 2013). Word2Vec has been shown to be effective for identifying similar test cases (Viggiato et al., 2022). In this study, we used the Continuous Bag-of-Words (CBOW) architecture with a window size of $10$ for Word2Vec, as this configuration yielded the highest FDR for TSM. This was determined through experiments evaluating both the CBOW and Skip-Gram architectures with window sizes ranging from 22 to $10$, in increments of 22. We trained Word2Vec models on the entire test case corpus and transformed each word into $300$-dimensional numerical vectors.
@@
GloVe (Pennington et al., 2014). GloVe is an unsupervised learning algorithm that leverages global word-word co-occurrence statistics from a corpus to learn word embeddings (Pennington et al., 2014). Its core idea is that the ratio of co-occurrence probabilities can capture meaningful linear substructures within the word vector space. We selected a window size of 22 for GloVe based on its performance in hyperparameter tuning experiments with window sizes ranging from 22 to $10$, in increments of 22. As in Word2Vec, we used GloVe to generate a $300$-dimensional numerical vector representation for each word in the test cases.
@@
FastText (Bojanowski et al., 2017). FastText is a word embedding model proposed by Bojanowski et al. (Bojanowski et al., 2017), which extends Word2Vec by incorporating subword information. Each word is represented as a combination of character n-grams, enabling the model to better handle rare and out-of-vocabulary words, and to generalize more effectively—particularly in morphologically rich or noisy text. We used the CBOW architecture with a window size of 88 as the FastText configuration, based on experiments evaluating both CBOW and Skip-Gram architectures with window sizes ranging from 22 to $10$ in increments of 22. We used FastText to convert each word in the test cases into a $300$-dimensional numerical vector.
$4.$ Baselines
@@
We compare RTM against (1) Random Minimization, a standard TSM baseline (Pan et al., 2023), (2) FAST-R approaches (Cruciani et al., 2019), a set of novel TSM approaches with two versions: fixed-size budget version and adequate version. The former satisfies the pre-defined minimization budget constraint, whereas the latter satisfies the full requirement coverage constraint. While being a widely used approach for TSM (Khan et al., 2018), Greedy-based TSM approaches iteratively select the test cases that maximize the coverage until the full coverage constraint is satisfied (Miranda and Bertolino, 2017; Noemmer and Haas, 2019). However, in our dataset, each test case covers only one requirement, rendering this option equivalent to Random Minimization with a requirement coverage constraint, as described below.
@@ 4.4.1. Random Minimization
Since the goal of TSM is to reduce the number of test cases and, consequently, lower testing costs, the most straightforward approach is to randomly remove test cases to meet the minimization budget (Ali et al., 2010). We evaluate the performance of RTM against Random Minimization with and without the requirement coverage constraint, denoted as RM-Req and RM-NoReq, respectively. For the version with the constraint, we applied the same selection strategy as GA initialization described in Section 3.4.1. For the version without the constraint, we randomly select $k$ test cases from the test suite, ensuring that $k$ adheres to the minimization budget.
4.4.2. FAST-R (Cruciani et al., 2019)
FAST-R is a family of similarity-based TSM approaches, including FAST++, FAST-CS, FAST-pw, and FAST-all. FAST++ and FAST-CS utilize k-means++ clustering (Arthur and Vassilvitskii, 2007) and constructed coresets (Bachem et al., 2018). In contrast, FAST-pw and FAST-all leverage minhashing and locality-sensitive hashing (Leskovec et al., 2014). FAST-R can be applied in two scenarios: the fixed-budget and adequate scenarios. The fixed-budget version of FAST-R minimizes the test suite within a predefined budget, while the adequate version prioritizes preserving the requirement coverage of the minimized test suite over adhering to the budget. One limitation of FAST-R is that it cannot simultaneously guarantee both a minimization budget and full requirement coverage. We compare RTM against FAST-R under both scenarios where FAST-R is applicable (i.e., fixed-budget and adequate test suite minimization). For FAST-R, we relied on the publicly available replication package [^3] https://github.com/ICSE19-FAST-R/FAST-R provided by its authors to ensure a fair and consistent comparison.
@@ 5.1. Performance of RTM under different configurations (RQ1)
Interestingly, despite being a simple, sparse representation, TF-IDF outperformed all other sentence-level embeddings. The characteristics of our industrial test cases can explain this. As shown in Figure 1, the test cases in our dataset follow structured and repetitive templates with little linguistic variation. The steps in the test cases differ only in variable names or parameter values, and TF-IDF precisely captures these differences through its word-frequency mechanism, down-weighting common words that appear in every test case and up-weighting rare words, including distinct variable names and parameter values. In contrast, other sentence-level embeddings, while capable of capturing deep semantic relationships, tend to smooth out subtle lexical differences that are crucial for distinguishing test cases in our dataset. This suggests that for template-based, domain-specific test descriptions, TF–IDF, which suppresses common terms and emphasizes rare tokens, often outperforms semantically rich embeddings that smooth over token-level distinctions.
@@
Answering RQ1: TF-IDF with cosine similarity yields the highest overall FDR (86.09%) under Preprocessing Method PM3 and Initialization Strategy 2 at the 50% minimization budget.
Table 1. FDR (%) of RTM across different configurations under 50% minimization budget.
@@ 5.2. Performance of RTM compared with baselines (RQ2)
We evaluated RTM against Random Minimization, which also employs Initialization Strategy 2, two versions of FAST-R (Cruciani et al., 2019)—the fixed-budget and adequate versions. For the fixed-budget version of FAST-R, we varied the minimization budget from $10\%$ to $90\%$, in increments of $10\%$. The fixed-budget version of FAST-R and RM-NoReq do not guarantee the 100% requirement coverage and are included in this RQ solely for FDR comparison. In the adequate scenario, we first ran FAST-R to determine the resulting minimization budget. This budget was then applied consistently to RTM and other baselines for experimental comparison. Notably, due to redundancy in our dataset, as few as $7\%$ (i.e., $54$ test cases, one test case per requirement) are sufficient to achieve $100\%$ requirement coverage.
@@
We observe that RTM consistently outperforms Random Minimization and FAST-R approaches in terms of FDR across various minimization budgets. The performance gap is particularly notable at mid-range budgets ($30\%$ to $60\%$). For instance, at a $40\%$ minimization budget, RTM achieves an FDR of $78.68\%$, whereas RM-Req and FAST++ yield approximately $65.45\%$ and $56.50\%$, respectively. This demonstrates RTM’s superior ability to retain fault-revealing test cases, especially under mid-range budget constraints. At lower and higher minimization budgets, the FDR of all techniques tends to converge toward either 0 or 11, thus making the difference in FDR between RTM and other techniques smaller. However, mid-range minimization budgets are more important in practice, as they offer more useful trade-offs between cost savings and fault detection.
@@
Table 2. FDR (%) and requirement coverage (%) comparison between RTM and baselines under the adequate-scenario setting.
@@
Figure 3. Comparison of FDR across minimization budgets for RTM, FAST-R, and Random Minimization, alongside the theoretical upper bound.
@@ 5.3. Impact of test suites redundancy (RQ3)
As expected, the FDR of all techniques improves as the test suite redundancy level increases, indicating that the TSM techniques tend to perform better–achieving higher FDR–when applied to more redundant test suites. This trend, however, becomes less evident at higher budgets ($80\%$ and $90\%$), where the FDR of all techniques converge toward 11. Random Minimization (RM-Req and RM-NoReq) slightly outperforms other FAST-R variants but consistently falls short of RTM. We also observed that across all minimization budgets and redundancy levels, RM-Req consistently outperformed RM-NoReq in FDR. This demonstrates that, for our dataset, explicitly ensuring that every requirement is covered not only preserves full requirement coverage but also significantly boosts the test suite’s ability to detect faults. In other words, prioritizing requirement coverage is an effective heuristic for improving fault detection under budgetary constraints.
@@
Figure 4. Comparison of FDR across varying redundancy levels for RTM and baseline approaches under seven minimization budgets.
@@ 5.4. Scalability of RTM compared with baselines (RQ4)
We observe that, although being slower than all other baselines, RTM minimizes test cases in seconds–$1.32$s to $4.44$s in terms of total minimization time–for test suite size from $187$ to $616$, which is computationally efficient. Moreover, the fact that the total minimization time of RTM follows a linear relationship with the test suite size indicates that RTM is scalable for larger test suites. According to the fitted linear regression model, when the test suite size $n$ is 200,000200{,}000, the predicted total minimization time of RTM is $6.217686\times 10^{-3}\,n+2.523669\approx 1246.06$ seconds, whereas FAST-CS (the fastest FAST-R approach) is $1.195291\times 10^{-4}\,n+0.051979\approx 23.96$ seconds and RM-Req is $9.062634\times 10^{-7}*n+0.000472\approx 0.18$ seconds. Thus, the time differences are $20.37$ mins and $20.76$ mins when compared to FAST-CS and RM-Req, respectively. Given the fact that, unlike test case selection and prioritization, test suite minimization only performs occasionally, such as major releases with many new added test cases and redundant old test cases (Pan et al., 2024, 2023), such differences are practically acceptable.
@@
In summary, RTM runs within 55 seconds across all evaluated test suites. Although being slower than the baselines, RTM scales approximately linearly with test suite sizes, and even for substantially larger suites, the absolute runtime gaps remain acceptable in practice—they are not expected to have practical consequences for TSM. Given that RTM consistently achieves much higher FDR across various datasets and minimization budgets than the baselines while preserving full requirement coverage, it is therefore a better option in many practical contexts.
@@
Figure 5. Scatter plots of the number of test cases and preparation time, search time and total minimization time (in sec) for RTM and baselines under 50% minimization budget.
@@ 7. Related Work
Several approaches have been proposed to support TSM, which can be classified in to three categories (Khan et al., 2018; Cruciani et al., 2019; Pan et al., 2023): greedy (Noemmer and Haas, 2019; Miranda and Bertolino, 2017), clustering (Coviello et al., 2018; Liu et al., 2011; Viggiato et al., 2022; Cruciani et al., 2019), and search-based (Zhang et al., 2019; Hemmati et al., 2013; Pan et al., 2023, 2024; Marchetto et al., 2017), plus hybrid combinations (Xia et al., 2021; Anwar et al., 2019; Yoo and Harman, 2010) thereof.
@@
Greedy-based approaches follow a heuristic, step-by-step selection strategy, where the test case that provides the most benefit toward the objective is chosen at each iteration, until the given constraints (e.g., full requirement coverage or code path coverage) are satisfied. Miranda et al. (Miranda and Bertolino, 2017) and Noemmer et al. (Noemmer and Haas, 2019) both employed greedy heuristics for TSM based on statement coverage. Their approaches iteratively selected test cases to maximize code coverage, achieving reasonable FDRs while significantly reducing the number of executed test cases. Greedy-based approaches are computationally efficient and easy to implement, but may suffer from suboptimal global performance due to their locally optimal decisions. Moreover, given that, in our dataset, each test case covers only one requirement, the coverage-based greedy strategy–iteratively selecting the test cases that maximize the coverage until the full coverage constraint is satisfied–is equivalent to Random Minimization with the requirement constraint (RM-Req), as described in Section $4.$
@@
Clustering-based approaches group similar test cases based on syntactic or semantic similarities, such as those derived from natural language descriptions or execution behavior. The goal is to reduce redundancy by selecting representative test cases from each cluster. These approaches are particularly useful in settings where test cases are written in natural language and code coverage information is unavailable. Viggiato et al. (Viggiato et al., 2022) proposed a comprehensive approach that combines text embedding, similarity measurement, and clustering to identify and remove similar test cases written in natural language. They split the requirement test cases into test steps and then calculate the similarity between test steps. Cruciani et al. (Cruciani et al., 2019) proposed FAST-R, a black-box TSM approach that leverages the source code of test cases without requiring execution or coverage information. FAST-R transforms test code into vector representations using a term frequency model, followed by dimensionality reduction via random projection. Similarly, Coviello et al. (Coviello et al., 2018) proposed a clustering-based approach using Hierarchical Agglomerative Clustering to perform TSM.
@@
Search-based approaches formulate the minimization problem as an optimization task and utilize metaheuristic algorithms (e.g., GA, simulated annealing) to explore the space of possible test subsets. These methods can balance multiple conflicting objectives, such as minimizing the number of test cases while maximizing fault detection capability and maintaining requirement coverage. Hemmati et al. (Hemmati et al., 2013) applied a search algorithm that minimizes test suites based on model-derived test case similarity. Similarly, Zhang et al.(Zhang et al., 2019) proposed UncerTest, a model-based test case minimization framework that leverages multi-objective search algorithms, including NSGA-II, MOCell(Nebro et al., 2009), and SPEA2 (Zitzler et al., 2001). ATM (Pan et al., 2023) is a black-box TSM approach that combines test code similarity analysis with evolutionary search. It represents test code using Abstract Syntax Trees and employs GA and Non-Dominated Sorting Genetic Algorithm II (Meyarivan et al., 2002) as the underlying evolutionary algorithms. Furthermore, LTM (Pan et al., 2024) leverages LLMs (CodeBERT, UniXcoder, and CodeLlama) to compute similarity between test cases and employs a GA to search for an optimal subset of test cases under a fixed budget. MORE+ (Marchetto et al., 2017) is a multi-objective TSM approach that uses NSGA-II to optimize structural, functional, and cost-related objectives, aiming to enhance fault detection while reducing redundancy and execution time. Although typically more computationally intensive than greedy and clustering approaches, search-based techniques support multiple minimization objectives and often achieve better performance for TSM.
@@
Hybrid approaches integrate complementary methods into a unified framework to balance multiple objectives, including reducing test suite size, preserving fault detection capability, and satisfying requirement or coverage constraints. Xia et al. (Xia et al., 2021) apply clustering-based test suite reduction combined with evolutionary multi-objective optimization. Anwar et al. (Anwar et al., 2019) combined a GA and particle swarm optimization to optimize regression test suites. Yoo et al. (Yoo and Harman, 2010) employ a hybrid multi-objective GA that integrates the efficient approximation of the greedy approach with the population-based GA’s ability to generate higher-quality Pareto fronts.
@@
While promising, existing techniques primarily focus on code-based solutions for minimization. In contrast, our work addresses minimization in the context of requirement-driven testing, where test cases are specified in natural language, which introduces distinct challenges. Moreover, current approaches do not explicitly support TSM under a fixed budget while simultaneously enforcing full requirement coverage as a hard constraint. To this end, we adopt a GA, as in the SOTA TSM work (i.e., ATM and LTM), and augment it with an explicit requirement coverage constraint that guarantees $100\%$ coverage under a fixed minimization budget. Note that we did not use NSGA-II, as reported in the ATM work, since it achieves comparable performance while being substantially more time-consuming than the GA. To the best of our knowledge, this is also the first work to investigate the impact of redundancy levels on fault detection capability under such constraints, providing valuable insights for future research and practical applications. Note that we retained FAST-R as one of the baselines, though it cannot simultaneously satisfy both the fixed-size minimization budget and the full coverage adequacy constraint. This is because FAST-R can be applied in two scenarios (i.e., fixed-size budget and adequate), which enables experimental comparison with our approach.